# UAV-Mounted IRS (UMI) in the Presence of Hovering Fluctuations: 3D Pattern Characterization and Performance Analysis

### v2

@@ Abstract
This paper investigates unmanned aerial vehicle (UAV)-mounted intelligent reflecting surfaces (IRS) to leverage the benefits of this technology for future communication networks, such as 6G. Key advantages include enhanced spectral and energy efficiency, expanded network coverage, and flexible deployment. One of the main challenges in employing UAV-mounted IRS (UMI) technology is the random fluctuations of hovering UAVs. Focusing on this challenge, this paper explores the capabilities of UMI with passive/active elements affected by UAV fluctuations in both horizontal and vertical angles, considering the three-dimensional (3D) radiation pattern of the IRS. The relationship between UAV fluctuations and IRS pattern is investigated by taking into account the random angular vibrations of UAVs. A tractable and closed-form distribution function for the IRS pattern is derived, using linear approximation and by dividing it into several sectors. In addition, closed-form expressions for outage probability (OP) are obtained using central limit theorem (CLT) and Gamma approximation. The theoretical expressions are validated through Monte Carlo simulations. The findings indicate that the random fluctuations of hovering UAVs have a notable impact on the performance of UMI systems. To avoid link interruptions due to UAV instability, IRS should utilize fewer elements, even though this leads to a decrease in directivity. As a result, unlike terrestrial IRS, incorporating more elements into aerial IRS systems does not necessarily improve performance due to the fluctuations in UAV. Numerical results show that the OP can be minimized by selecting the optimal number of IRS elements and using active elements.
Index Terms:
Intelligent reflecting surfaces (IRS), IRS 3D pattern, unmanned aerial vehicle (UAV), UAV fluctuations, UAV-mounted IRS (UMI).
@@ I Introduction
The intelligent reflecting surface (IRS) is a promising technology suggested for future communication networks. In recent years, this technology has gained significant attention due to its capability to manage the wireless environment between transceivers[1]. Specifically, an IRS consists of numerous passive/active elements that act as smart reflectors, reflecting the incoming signals in desired direction by adjusting their phase shifts with a programmable controller. As a result, the signals reflected by individual IRS elements can be combined with multipath signals, thereby amplifying the received signal strength and reducing interference. This process contributes to a notable improvement in spectral efficiency (SE) [2]. Unlike conventional active relays like decode-and-forward (DF) and amplify-and-forward (AF) relays, the passive IRS is composed of passive components and does not require RF chains. Furthermore, IRS reflects incoming signals with minimal power usage, leading to low energy consumption [3]. Because of their straightforward installation process and flexible shape, IRS can be installed on building facades or directly onto drones to help terrestrial user equipment (UEs) with poor connectivity [4, 5]. The features mentioned above make IRS not only a promising solution for green networks but also a valuable complement to other existing technologies. However, the performance of passive IRS-assisted systems may be restricted by significant product-distance path loss. One potential solution to this challenge involves increasing the number of passive reflecting elements to leverage the square-order beamforming gain. Alternatively, deploying a passive IRS in close proximity to the transmitter and/or receiver can effectively mitigate path loss [6]. On the other hand, a novel variant of IRS, known as active IRS, has recently been introduced in [7, 8, 9, 10]. This approach addresses the drawbacks of passive IRS by employing low-cost components to amplify the reflected signals. Typically, an active IRS comprises several active reflecting elements that integrate negative resistance components like tunnel diodes and negative impedance converters. This allows the IRS to reflect incoming signals while boosting their power [11]. In contrast to AF relays, which depend on power-hungry RF chains, active IRS reflects signals directly through low-power reflection-type amplifiers.
@@ I-A Related Works
Beyond separate IRS-UAV systems, recent research has introduced UAV-mounted IRS (UMI) configurations, where the IRS is attached to or carried by the UAV. For instance, [19] proposes a communication scheme where an IRS is mounted on a UAV to facilitate connections between a base station (BS) and ground users. This study introduces methods to enhance both SE and EE by jointly optimizing active beamforming, passive beamforming, and UAV trajectory. In [20], a UAV carrying an IRS is proposed as a relay node to support covert communication systems (CCS). To ensure covertness, the study calculates the minimum error detection probability at the eavesdropper and optimizes the UAV trajectory and IRS phase shifts to maximize the average communication rate. In [21], the authors explore the security of simultaneous wireless information and power transfer (SWIPT) systems supported by IRS and UAVs, taking into account the energy consumption of rotary-wing UAVs during flight. In this setup, an IRS mounted on a UAV improves the quality of legitimate transmissions, while artificial noise generated at the base station (BS) is used to disrupt eavesdropping attempts. Ground devices (GDs) utilize power splitting (PS) technology to concurrently decode information and harvest energy. The study focuses on optimizing BS transmit beamforming, UAV-IRS phase shifts, trajectory/velocity, and GD PS ratios to maximize the overall secrecy rate for all GDs.
@@
Note that the hovering fluctuations of UAVs are not considered in the studies mentioned above. While the UAV jitter is studied in [22, 23, 24], characterizing the 3D IRS pattern and its distribution under UAV fluctuations remains unattended. The authors in [22] analyze the use of a multi-aerial IRS in a secure SWIPT system, considering UAV jitter. Angle estimation errors caused by UAV jitter are converted into CSI errors using linear approximation techniques. Subsequently, a joint optimization problem is formulated to maximize the average secrecy rate, incorporating the beamforming vector, IRS phase shift matrices, and UAV trajectories. Random airflow and fuselage vibrations can greatly affect the communication capabilities of UAVs. In [23], a novel active IRS is introduced to address this issue, enabling a secure and energy-efficient beamforming design. The proposed framework accounts for the effects of UAV jitter and involves the joint optimization of the active IRS reflection coefficient, UAV-based BS beamforming, and UAV trajectory, while ensuring compliance with worst-case secrecy rate constraints. In [24], the authors investigate a UAV-assisted multi-user IRS communication system, aiming to minimize power consumption through the joint optimization of active beamforming, passive beamforming, and UAV trajectories. The study takes into account practical constraints such as UAV jitters and imperfections in hardware components.
@@ I-B Our Contribution
Considering the current state of above research, this study examines a UMI system equipped with both passive and active elements, taking into account the effects of UAV fluctuations. Due to the UAV instability, we assess how its horizontal and vertical angle fluctuations affect system performance. The main contribution of this paper is the development of a 3D IRS pattern, along with the determination of its distribution under UAV fluctuations, and the analysis of the outage probability (OP) performance of the UMI. By leveraging the derived analytical expressions, we minimize the OP under different levels of UAV fluctuations. This is achieved by selecting an optimal number of IRS elements, which ultimately enhances system reliability. The main contributions of this work are outlined as follows
1) We address the issue of instability in real-world UAV platforms, which are susceptible to airflow disturbances and airframe vibrations, leading to horizontal and vertical angular fluctuations. To conduct performance analysis, it is necessary to have the distribution of IRS 3D pattern. Therefore, we characterize the 3D pattern of the IRS and determine its distribution while considering UAV fluctuations. Note that this distribution is applicable in all systems and channel models involving UMI for performance analysis under UAV fluctuations.
2) In order to determine the IRS 3D pattern distribution which is a product of the array factor and single radiation pattern, we follow a three-step analytical approach. First, we obtain the distribution of the array factor then, we obtain the distribution of the single radiation pattern. Finally, we derive the tractable and closed-form probability distribution function (PDF) of the IRS 3D pattern by utilizing the results of the previous two steps. To simplify the analysis, we apply linear approximations to address the error variations in elevation and azimuth angles caused by UAV fluctuations. Furthermore, we introduce a sectoral model for the array factor and single radiation pattern, resulting in a more straightforward approach for determining the distribution of the IRS 3D pattern.
3) Using the derived PDF of the IRS 3D pattern, we establish closed-form expressions for the OP of the UMI in four scenarios: single-input single-output (SISO)-Passive-IRS, multiple-input single-output (MISO)-Passive-IRS, SISO-Active-IRS, and MISO-Active-IRS. For this analysis, we apply the central limit theorem (CLT) and Gamma approximation techniques.
@@ I-C Outline and Notations
Notation: Throughout the paper, matrices and column vectors are denoted by bold uppercase and bold lowercase letters, respectively. The transpose and Hermitian transpose of a vector or matrix are indicated by (.$(.)^{T}$ and (.$(.)^{H}$, respectively. $\mathbf{0}_{N}$ denotes the zero vector of size $N$, where every component is 00. The $N\times N$ identity matrix is expressed by $\mathbf{I}_{N}$. The Kronecker product of two vectors is denoted by $\otimes$. A real and complex normal distribution with mean $\mu$ and variance $\sigma^{2}$ is expressed as $\mathcal{N}(\mu,\sigma^{2})$ and $\mathcal{CN}(\mu,\sigma^{2})$, respectively. The Q-function is denoted by $Q(.)$, while the Dirac delta function is represented by $\delta(.)$. Finally, the expectation operation is expressed as $\mathbb{E}(.)$.
@@ II System Model
Without loss of generality, a 3D Cartesian coordinate system is adopted. The positions of the BS, IRS, and UE are defined as $\mathbf{u}_{\footnotesize\text{BS}}=[x_{\footnotesize\text{BS}},y_{\footnotesize\text{BS}},h_{\footnotesize\text{BS}}]$, $\mathbf{u}_{\footnotesize\text{IRS}}=[x_{\footnotesize\text{IRS}},y_{\footnotesize\text{IRS}},h_{\footnotesize\text{IRS}}]$, and $\mathbf{u}_{\footnotesize\text{UE}}=[x_{\footnotesize\text{UE}},y_{\footnotesize\text{UE}},0]$, respectively. The transmitted signal to IRS and reflected signal from IRS is represented with the subscript $q\in\{t,r\}$ where $t$ and $r$ indicate the transmitted signal and the reflected signal, respectively. As shown in Fig. 2, $\theta_{q}$ and $\phi_{q}$ denote the elevation and azimuth angles from the IRS in the direction of the BS/UE. Assume $\varepsilon_{x}$ and $\varepsilon_{y}$ are the angular variations of UAV along the $x-z$ and $y-z$ axes, respectively. According to the central limit theorem, the fluctuations in the orientations of the UAV have Gaussian distribution [25, 26, 27]. That is, we have $\varepsilon_{x}\sim\mathcal{N}(\mu_{x},\sigma^{2}_{x})$, and $\varepsilon_{y}\sim\mathcal{N}(\mu_{y},\sigma^{2}_{y})$.
@@
We assume the IRS is configured as a uniform square array with $N=N_{q}\times N_{q}$ elements. To ensure independence between IRS elements, the spacing between elements in both $x-$ and $y-$directions, denoted as $d_{x}$ and $d_{y}$, respectively, are set to half of the wavelength, $\frac{\lambda}{2}$. By applying the basic trigonometric formulas, the angles of elevation and azimuth from the IRS towards the direction of the BS/UE without any UAV fluctuations, denoted as $\theta_{q}$ and $\phi_{q}$ are obtained as
$\theta_{q}=\arctan(\sqrt{\tan^{2}(\theta_{qx})+\tan^{2}(\theta_{qy})}),$ (1)
$\phi_{q}=\arctan(\frac{\tan(\theta_{qy})}{\tan(\theta_{qx})}),$
@@
where the directions of IRS pattern without fluctuations are represented by $\theta_{qx}$ and $\theta_{qy}$ in the $x-z$ and $y-z$ Cartesian coordinates, respectively. On the other hand, the elevation and azimuth angles under UAV fluctuaions, i.e., $\tilde{\theta}_{q}$ and $\tilde{\phi}_{q}$, are calculated as
$\tilde{\theta}_{q}(\varepsilon_{x},\varepsilon_{y})=\tan^{-1}(\sqrt{\tan^{2}(\theta_{qx}+\varepsilon_{x})+\tan^{2}(\theta_{qy}+\varepsilon_{y})}),$ (2)
$\tilde{\phi}_{q}(\varepsilon_{x},\varepsilon_{y})=\tan^{-1}(\frac{\tan(\theta_{qy}+\varepsilon_{y})}{\tan(\theta_{qx}+\varepsilon_{x})}).$

### v3

@@ Abstract
This paper investigates unmanned aerial vehicle (UAV)-mounted intelligent reflecting surfaces (IRS) to leverage the benefits of this technology for future communication networks, such as 6G. Key advantages include enhanced spectral and energy efficiency, expanded network coverage, and flexible deployment. One of the main challenges in employing UAV-mounted IRS (UMI) technology is the random fluctuations of hovering UAVs. Focusing on this challenge, this paper explores the capabilities of UMI with passive/active elements affected by UAV fluctuations in both horizontal and vertical angles, considering the three-dimensional (3D) radiation pattern of the IRS. The relationship between UAV fluctuations and IRS pattern is investigated by taking into account the random angular vibrations of UAVs. A tractable and closed-form distribution function for the IRS pattern is derived, using linear approximation and by dividing it into several sectors. In addition, closed-form expressions for outage probability (OP) are obtained using the central limit theorem (CLT) and the Gamma approximation. The theoretical expressions are validated through Monte Carlo simulations. Our findings indicate that the random fluctuations of hovering UAVs have a notable impact on the performance of UMI systems. To avoid link interruptions due to UAV instability, IRS should utilize fewer elements, even though this leads to a decrease in directivity. As a result, unlike terrestrial IRS, incorporating more elements into aerial IRS systems does not necessarily improve performance due to the fluctuations in UAVs. Numerical results show that the OP can be minimized by selecting the optimal number of IRS elements and using active elements.
@@ I Introduction
The intelligent reflecting surface (IRS) is a promising technology suggested for future communication networks. In recent years, this technology has gained significant attention due to its capability to manage the wireless environment between transceivers [1, 2]. Specifically, an IRS consists of numerous passive/active elements that act as smart reflectors, reflecting the incoming signals in the desired direction by adjusting their phase shifts with a programmable controller. As a result, the signals reflected by individual IRS elements can be combined with multipath signals, thereby amplifying the received signal strength and reducing interference. This process contributes to a notable improvement in spectral efficiency (SE) [3]. Unlike conventional active relays like decode-and-forward (DF) and amplify-and-forward (AF) relays, the passive IRS is composed of passive components and does not require RF chains. Furthermore, IRS reflects incoming signals with minimal power usage, leading to low energy consumption [4]. Because of their straightforward installation process and flexible shape, IRS can be installed on building facades or directly onto drones to help terrestrial user equipment (UEs) with poor connectivity [5, 6]. The features mentioned above make IRS not only a promising solution for green networks but also a valuable complement to other existing technologies. However, the performance of passive IRS-assisted systems may be restricted by significant product-distance path loss. One potential solution to this challenge involves increasing the number of passive reflecting elements to leverage the square-order beamforming gain. Alternatively, deploying a passive IRS in proximity to the transmitter and/or receiver can effectively mitigate path loss [7]. On the other hand, a novel variant of IRS, known as active IRS, has recently been introduced in [8, 9, 10, 11]. This approach addresses the drawbacks of passive IRS by employing low-cost components to amplify the reflected signals. Typically, an active IRS comprises several active reflecting elements that integrate negative resistance components like tunnel diodes and negative impedance converters. This allows the IRS to reflect incoming signals while boosting its power [12]. In contrast to AF relays, which depend on power-hungry RF chains, active IRS reflects signals directly through low-power reflection-type amplifiers.
@@ I-A Related Works
Beyond separate IRS-UAV systems, recent research has introduced UAV-mounted IRS (UMI) configurations, where the IRS is attached to or carried by the UAV. The key advantage of the UMI is its significantly lower power consumption compared to active relays. This is because the UMI reflects signals without needing power-hungry components, requiring only minimal energy for its phase-shift controller [26, 27]. While this power draw is small relative to the UAV’s propulsion, it is still a factor; thus, mission time can be extended using energy-efficient control algorithms and renewable energy sources like solar panels [28]. In [20], a communication scheme is proposed where an IRS is mounted on a UAV to facilitate connections between a base station (BS) and ground users. This study introduces methods to enhance both SE and EE by jointly optimizing active beamforming, passive beamforming, and UAV trajectory. In [21], a UAV carrying an IRS is proposed as a relay node to support covert communication systems (CCS). To ensure covertness, the study calculates the minimum error detection probability at the eavesdropper and optimizes the UAV trajectory and IRS phase shifts to maximize the average communication rate. In [22], the authors explore the security of simultaneous wireless information and power transfer (SWIPT) systems supported by IRS and UAVs, taking into account the energy consumption of rotary-wing UAVs during flight. In this setup, an IRS mounted on a UAV improves the quality of legitimate transmissions, while artificial noise generated at the base station (BS) is used to disrupt eavesdropping attempts. Ground devices (GDs) utilize power splitting (PS) technology to concurrently decode information and harvest energy. The study focuses on optimizing BS transmit beamforming, UAV-IRS phase shifts, trajectory/velocity, and GD PS ratios to maximize the overall secrecy rate for all GDs.
@@
Recently, UAV jitter in the presence of IRS is studied in [23, 24, 25]. The authors in [23] analyze the use of a multi-aerial IRS in a secure SWIPT system, considering UAV jitter. Angle estimation errors caused by UAV jitter are converted into CSI errors using linear approximation techniques. Subsequently, a joint optimization problem is formulated to maximize the average secrecy rate, incorporating the beamforming vector, IRS phase shift matrices, and UAV trajectories. Random airflow and fuselage vibrations can greatly affect the communication capabilities of UAVs. In [24], a novel active IRS is introduced to address this issue, enabling a secure and energy-efficient beamforming design. The proposed framework accounts for the effects of UAV jitter and involves the joint optimization of the active IRS reflection coefficient, UAV-based BS beamforming, and UAV trajectory, while ensuring compliance with worst-case secrecy rate constraints. In [25], the authors investigate a UAV-assisted multi-user IRS communication system, aiming to minimize power consumption through the joint optimization of active beamforming, passive beamforming, and UAV trajectories. The study takes into account practical constraints such as UAV jitters and imperfections in hardware components.
@@ I-B Our Contribution
While prior studies have investigated IRS-assisted UAV communications, the impact of UAV hovering fluctuations remains largely unaddressed [15, 16, 17, 18, 19, 20, 21, 22]. Although recent works have begun to explore UAV jitter in IRS-assisted scenarios [23, 24, 25], critical gaps persist. For instance, some prior studies formulate the problem as an optimization of beamforming, IRS phase shifts, and other parameters, yet fail to provide closed-form derivations. Moreover, while the number of IRS elements should be optimized under UAV fluctuations, existing work typically assumes a fixed number and lacks analytical derivations to determine the optimal configuration. To address these limitations, this work provides a comprehensive characterization of the 3D IRS pattern under UAV hovering fluctuations, analyzes its statistical distribution, and derives the optimal number of elements for both passive and active IRS architectures. By incorporating both IRS types and explicitly modeling UAV fluctuation effects, our analysis provides new theoretical insights into UMI system design. Due to the UAV’s instability, we assess how its horizontal and vertical angle fluctuations affect system performance. The main contribution of this paper is the development of a 3D IRS pattern, along with the determination of its distribution under UAV fluctuations, and the analysis of the outage probability (OP) of the UMI. By leveraging the derived analytical expressions, we minimize the OP under different levels of UAV fluctuations. This is achieved by selecting an optimal number of IRS elements, which ultimately enhances system reliability. The main contributions of this work are outlined as follows
1) We address the issue of instability in real-world UAV platforms, which are susceptible to airflow disturbances and airframe vibrations, leading to horizontal and vertical angular fluctuations. To conduct performance analysis, it is necessary to have the distribution of the IRS 3D pattern. Therefore, we characterize the 3D pattern of the IRS and determine its distribution while considering UAV fluctuations. Note that this distribution is applicable in all systems involving UMI for performance analysis under UAV fluctuations.
2) To determine the IRS 3D pattern distribution, which is a product of the array factor and single radiation pattern, we follow a three-step analytical approach. First, we obtain the distribution of the array factor, then we treat the single radiation pattern as deterministic. Finally, we derive the tractable and closed-form probability distribution function (PDF) of the IRS 3D pattern by utilizing the results of the previous two steps. To simplify the analysis, we apply linear approximations to address the error variations in elevation and azimuth angles caused by UAV fluctuations. Furthermore, we introduce a sectoral model for the array factor and single radiation pattern, resulting in a more straightforward approach for determining the distribution of the IRS 3D pattern.
3) Using the derived PDF of the IRS 3D pattern, we establish closed-form expressions for the OP of the UMI in two scenarios: Passive-IRS and Active-IRS. For this analysis, we apply the central limit theorem (CLT) and Gamma approximation techniques.
@@ I-C Outline and Notations
Notation: Throughout the paper, matrices and column vectors are denoted by bold uppercase and bold lowercase letters, respectively. The transpose and Hermitian transpose of a vector or matrix are indicated by $(.)^{T}$ and $(.)^{H}$, respectively. $\mathbf{0}_{N}$ denotes the zero vector of size $N$, where every component is 0. The $N\times N$ identity matrix is expressed by $\mathbf{I}_{N}$. The Hadamard product and Kronecker product of two vectors are denoted by $\odot$ and $\otimes$, respectively. A real and complex normal distribution with mean $\mu$ and variance $\sigma^{2}$ is expressed as $\mathcal{N}(\mu,\sigma^{2})$ and $\mathcal{CN}(\mu,\sigma^{2})$, respectively. The Q-function is denoted by $Q(.)$, while the Dirac delta function is represented by $\delta(.)$. Finally, the expectation operation is expressed as $\mathbb{E}(.)$.
@@ II System Model
Without loss of generality, a 3D Cartesian coordinate system is adopted. The positions of the BS, IRS, and UE are defined as $\mathbf{u}_{\footnotesize\text{BS}}=[x_{\footnotesize\text{BS}},y_{\footnotesize\text{BS}},h_{\footnotesize\text{BS}}]$, $\mathbf{u}_{\footnotesize\text{IRS}}=[x_{\footnotesize\text{IRS}},y_{\footnotesize\text{IRS}},h_{\footnotesize\text{IRS}}]$, and $\mathbf{u}_{\footnotesize\text{UE}}=[x_{\footnotesize\text{UE}},y_{\footnotesize\text{UE}},0]$, respectively. The transmitted signal to IRS and reflected signal from IRS are represented with the subscript $q\in\{t,r\}$ where $t$ and $r$ indicate the transmitted signal and the reflected signal, respectively. As shown in Fig. 2, $\theta_{q}$ and $\phi_{q}$ denote the elevation and azimuth angles from the IRS in the direction of the BS/UE. Assume $\varepsilon_{x}$ and $\varepsilon_{y}$ are the angular variations of UAV along the xzx-z and yzy-z axes, respectively. According to the CLT, the fluctuations in the orientations of the UAV have a Gaussian distribution [29, 30, 31]. That is, we have $\varepsilon_{x}\sim\mathcal{N}(\mu_{x},\sigma^{2}_{x})$, and $\varepsilon_{y}\sim\mathcal{N}(\mu_{y},\sigma^{2}_{y})$.
@@ II-A Channel Model
By applying the basic trigonometric formulas, the angles of elevation and azimuth from the IRS towards the direction of the BS/UE without any UAV fluctuations, denoted as $\theta_{q}$ and $\phi_{q}$ are obtained as
$\theta_{q}=\arctan(\sqrt{\tan^{2}(\theta_{qx})+\tan^{2}(\theta_{qy})}),$ (1a)
$\phi_{q}=\arctan(\frac{\tan(\theta_{qy})}{\tan(\theta_{qx})}),$ (1b)
@@
where the directions of the IRS pattern without fluctuations are represented by $\theta_{qx}$ and $\theta_{qy}$ in the xzx-z and yzy-z Cartesian coordinates, respectively. On the other hand, the elevation and azimuth angles under UAV fluctuations, i.e., $\tilde{\theta}_{q}$ and $\tilde{\phi}_{q}$, are calculated as
$\tilde{\theta}_{q}(\varepsilon_{x},\varepsilon_{y})=\tan^{-1}(\sqrt{\tan^{2}(\theta_{qx}+\varepsilon_{x})+\tan^{2}(\theta_{qy}+\varepsilon_{y})}),$ (2a)
$\tilde{\phi}_{q}(\varepsilon_{x},\varepsilon_{y})=\tan^{-1}(\frac{\tan(\theta_{qy}+\varepsilon_{y})}{\tan(\theta_{qx}+\varepsilon_{x})}).$ (2b)
where depend on variables $\varepsilon_{x}$ and $\varepsilon_{y}$.
Since UMI is usually at altitudes high enough to establish LoS links with ground devices, it experiences low small-scale fading. Consequently, all channels–between BS and IRS, as well as between IRS and UE–are modeled as the Rician distribution [34], denoted by $\mathbf{H}_{BI}\in\mathbb{C}^{N\times M}$ and $\mathbf{h}_{IU}\in\mathbb{C}^{N\times 1}$, respectively. The array response of the BS is expressed as $\mathbf{a}_{B}(\theta,\phi)=[1,e^{-j\frac{2\pi}{\lambda}d\sin(\theta)\cos(\phi)},...,e^{-j\frac{2\pi}{\lambda}d(M-1)\sin(\theta)\cos(\phi)}]$ where $d$ is the adjacent antenna distance. The angles of elevation and azimuth from BS towards the direction of the IRS are denoted by $\theta_{0}$ and $\phi_{0}$, respectively. The array response of the IRS is given by
$\mathbf{a}_{I}(\theta,\phi)=[\mathbf{a}_{Ix}(\theta,\phi)\otimes\mathbf{a}_{Iy}(\theta,\phi)]^{T}.$ (3)
where
$\mathbf{a}_{Ix}(\theta,\phi)=[1,e^{-j\frac{2\pi}{\lambda}d_{x}\sin(\theta)\cos(\phi)},...,e^{-j\frac{2\pi}{\lambda}d_{x}(N_{q}-1)\sin(\theta)\cos(\phi)}],$ (4a)
$\mathbf{a}_{Iy}(\theta,\phi)=[1,e^{-j\frac{2\pi}{\lambda}dy\sin(\theta)\sin(\phi)},...,e^{-j\frac{2\pi}{\lambda}d_{y}(N_{q}-1)\sin(\theta)\sin(\phi)}].$ (4b)
Therefore, the channel of the BS-IRS and IRS-UE can be respectively given by
$\bar{\mathbf{H}}_{BI}=\sqrt{\beta_{0}E(\zeta_{t})}(\mathbf{H}_{BI}\odot(\mathbf{a}_{I}(\zeta_{t})\mathbf{a}_{B}(\theta_{0},\phi_{0}))),$ (5a)
$\bar{\mathbf{h}}_{IU}=\sqrt{\beta_{1}E(\zeta_{r})}(\mathbf{h}_{IU}\odot\mathbf{a}_{I}(\zeta_{r})),$ (5b)
where, $\zeta_{t}=\{\tilde{\theta}_{t},\tilde{\phi}_{t}\}$, $\zeta_{r}=\{\tilde{\theta}_{r},\tilde{\phi}_{r}\}$ and $E(\zeta)$ is calculated for $\zeta=\{\theta,\phi\}$ as
$E(\zeta)=\begin{cases}\cos^{3}(\theta),&\theta\in[0,\pi/2],\phi\in[0,2\pi]\\ 0,&\theta\in(\pi/2,\pi],\phi\in[0,2\pi].\end{cases}$ (6)
$\beta_{0}$ and $\beta_{1}$ are defined as $\beta_{k}\triangleq c_{0}{d_{k}}^{-\alpha_{k}},\quad k\in\{0,1\}$ where $c_{0}$ represents the reference path loss at a distance of 11m. The path loss exponents for BS-IRS and IRS-UE links are denoted by $\alpha_{0}$ and $\alpha_{1}$, respectively, whereas $d_{0}$ and $d_{1}$ represent the distance from the IRS to the BS and UE, respectively. The Rician factors of $\mathbf{H}_{BI}$ and $\mathbf{h}_{IU}$ are $K_{0}$ and $K_{1}$, respectively. We assume the small-scale fading is constant over a channel coherence interval (CCI), which is determined by both multi-path fading and UAV fluctuations. Since UAV fluctuations (e.g., wind or mechanical vibrations) occur on a much larger time scale, their variation within a single CCI is negligible.
II-B Received Signal Model
The reflection matrix of the active IRS is defined as $\mathbf{\Theta}\triangleq\mathbf{A}\mathbf{\Phi}$, where $\mathbf{\Phi}\triangleq\text{diag}(e^{j\phi_{1}},...,e^{j\phi_{N}})$ represents the IRS phase shift matrix with $\phi_{n}$ being the phase shift at the $n$-th element, $n\in\{1,2,...,N\}$. Additionally, the active IRS amplification matrix is denoted by $\mathbf{A}\triangleq\text{diag}(a_{1},...,a_{N})$ where $a_{n}$ is the amplification factor of the $n$-th element. Unlike passive IRS, active IRS introduces non-negligible thermal noise at each reflector, denoted as $\mathbf{n}_{F}\in\mathbb{C}^{N\times 1}$, which is typically modeled as $\mathbf{n}_{F}\sim\mathcal{N}(\mathbf{0}_{N},\sigma_{f}^{2}\mathbf{I}_{N})$, where $\sigma_{f}^{2}$ is the amplification noise power [11]. Finally, the received signal at the UE is expressed as
$y={\bar{\mathbf{h}}}^{H}_{IU}\mathbf{\Theta}{\bar{\mathbf{H}}}_{BI}\mathbf{w}s+{\bar{\mathbf{h}}}^{H}_{IU}\mathbf{\Theta}\mathbf{n}_{F}+n,$ (7)
where $n\sim\mathcal{N}(0,\sigma_{n}^{2})$ is the additive white Gaussian noise (AWGN) at the UE. The transmit signal from BS is represented by $s$ with transmit power $P_{t}$. The beamforming vector is denoted by $\mathbf{w}\in\mathbb{C}^{M\times 1}$. The power constraint at active IRS is expressed as follows [35]
$P_{t}\|\mathbf{A\Phi}{\bar{\mathbf{H}}}_{BI}\mathbf{w}\|^{2}+\|\mathbf{A\Phi}\mathbf{I}_{N}\|^{2}\sigma_{f}^{2}\leq P_{F},$ (8)
where $P_{F}$ is the maximum amplification power of the active IRS.
Since the CSI is sensitive to unforeseen feedback delays and IRS effects, the resulting residual CSI error can significantly degrade system performance [36, 37, 38]. Given this imperfect CSI at the transmitter, denoted as $\mathbf{h}$, its relationship with the perfect CSI can be modeled as follows [38]
$\hat{\mathbf{h}}_{k}=\sqrt{1-\zeta}\mathbf{h}+\sqrt{\zeta}\mathbf{h}_{e},$ (9)
where $\mathbf{h}^{H}={\bar{\mathbf{h}}}^{H}_{IU}\mathbf{\Theta}{\bar{\mathbf{H}}}_{BI}$ denotes the perfect cascaded channel, and $\mathbf{h}_{e}$ is the residual CSI error with zero-mean complex Gaussian entries of variance $\sigma_{e}^{2}=E\{|{h_{m}}|^{2}\}$, and is independent of $\mathbf{h}$. The CSI error factor $\zeta$, which ranges from 0 to $1$, indicates the accuracy of the CSI. Accounting for this imperfection, the received signal at the UE is written as
$\hat{y}=\sqrt{1-\zeta}\mathbf{h}^{H}\mathbf{w}s+\sqrt{\zeta}\mathbf{h}^{H}_{e}\mathbf{w}s+{\bar{\mathbf{h}}}^{H}_{IU}\mathbf{\Theta}\mathbf{n}_{F}+n.$ (10)
To provide a conservative performance bound ensuring reliable system design, we model the CSI error effect by incorporating its variance in the effective noise power $\sigma_{e}^{2}$. Therefore, the instantaneous SNR at the UE can be expressed as
$\gamma=\frac{P_{t}(1-\zeta)|{\bar{\mathbf{h}}}^{H}_{IU}\mathbf{A\Phi}{\bar{\mathbf{H}}}_{BI}\mathbf{w}|^{2}}{\|\bar{\mathbf{h}}^{H}_{IU}\mathbf{A\Phi}\|^{2}\sigma_{f}^{2}+P_{t}\zeta\|\mathbf{w}\|^{2}\sigma_{e}^{2}+\sigma_{n}^{2}}.$ (11)
As discussed in [39] and [40], we can assume a strong LoS component and negligible NLoS components for the BS-IRS link. This is due to the strong directivity of the multiple antennas at the BS and the low distance between the BS and the IRS. Based on the discussion in [35], for simplicity, we assume that each reflecting element has the same amplification factor $A$. To achieve maximum SNR, $\mathbf{w}^{*}$ is determined by the maximum ratio transmission (MRT) principle as [39]
$\mathbf{w}^{*}=\frac{(\bar{\mathbf{h}}^{H}_{IU}\mathbf{A}^{*}\mathbf{\Phi}^{*}\bar{\mathbf{H}}_{BI})^{H}}{\|\bar{\mathbf{h}}^{H}_{IU}\mathbf{A}^{*}\mathbf{\Phi}^{*}\bar{\mathbf{H}}_{BI}\|},$ (12)
which satisfies the unit power constraint $\|\mathbf{w}^{*}\|^{2}=1$. The optimal active-IRS reflection design is
$\mathbf{\Phi}^{*}]_{n}=e^{-j(\angle[{\mathbf{H}}_{BI}]_{m,n}+\angle[{\mathbf{h}}_{IU}^{H}]_{n}+[\mathbf{a}_{I}(\theta_{t},\phi_{t})]_{n}+[\mathbf{a}_{I}(\theta_{r},\phi_{r})]_{n})},~\forall n,$ (13)
$(A^{*})^{2}=\frac{P_{F}}{P_{t}\beta_{0}E(\zeta_{t})\sum_{n=1}^{N}|H_{BI,m,n}|^{2}+N\sigma_{f}^{2}}.$ (14)
Hence, the instantaneous maximum SNR can be expressed as
$\gamma_{\text{max}}=\frac{P_{t}M(1-\zeta)\beta_{0}\beta_{1}(A^{*})^{2}\mathbb{G}(\sum_{n=1}^{N}|H_{BI,m,n}||h_{IU,n}|)^{2}}{\beta_{1}(A^{*})^{2}E(\zeta_{r})\sigma_{f}^{2}\sum_{n=1}^{N}|h_{IU,n}|^{2}+P_{t}\zeta\sigma_{e}^{2}+\sigma_{n}^{2}},$ (15)